OpenStudy (anonymous):

Let f(x) = 24/3x -2 = . Find f(-2). I think it's -3. 3(-2) -2 = -8 24/8 = -3 Am I right?

OpenStudy (alexandervonhumboldt2):

look: substitute -2 for x: \[\frac{ 24 }{ 3 }*(-2)-2=8*(-2)-2=-16-2=-18\]

OpenStudy (anonymous):

Can't be that, my choices are: -6 -3 3 6 I chose -3.

OpenStudy (alexandervonhumboldt2):

oh i understood

OpenStudy (alexandervonhumboldt2):

your equation is 24/(3x-2)

OpenStudy (anonymous):

\[\frac{ 24 }{ 3x-2 }\] Is the problem written like this?

OpenStudy (alexandervonhumboldt2):

24/(3x-2)=24/(-6-2)=24/-8=-3 correct then

OpenStudy (anonymous):

Yes RKL, it is.

OpenStudy (alexandervonhumboldt2):

what you wrote @Tatsurion looked like \[\frac{ 24 }{ 3 }*x-2\]

OpenStudy (anonymous):

Thank you. I'm having a lot of function difficulties and I need to make sure I get things right.

OpenStudy (anonymous):

Okay then the answer will be 6.

OpenStudy (anonymous):

How is it 6? I see no way that could happen unless you did the negative incorrectly.

OpenStudy (anonymous):

If it as somehow a positive 4, then it would be 6.

OpenStudy (anonymous):

\[\frac{ 24 }{ 3(2)-2 } = \frac{ 24 }{ 4} = 6\]

OpenStudy (alexandervonhumboldt2):

@Rkl36097 the function is -2 not 2

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Hence why I said he did the negative wrong.

OpenStudy (anonymous):

Or she, I may be wrong.

OpenStudy (anonymous):

I am a guy and yes you guys are right I forgot about the negative. My bad.

OpenStudy (anonymous):

No problem!